OpenStudy (anonymous):

can someone help me with this. Let f(x) = x2 - 8x + 5. Find f(-1). ?

OpenStudy (anonymous):

1+8+5=14

OpenStudy (anonymous):

14 f(-1)=(-1)^2-8(-1)+5 =1+8+5 =14

OpenStudy (anonymous):

thanks